What is the mass in gram of 4.65*10^22 moles of P4O10
How many atoms are present in this sample
1
Expert's answer
2020-02-04T06:35:45-0500

1 atom= 6.022*10^ 23moles. Therefore,

4.65*10^ 22 moles have ;4.6*10 ^22moles*1atom/6.022*10^ 23moles=0.0764atoms

Mass of P4010 =molar mass*number of atoms

Mass = 0.0764*283.886g/mol=21.68g